The random variable Z has a standard normal distribution. Find the value of b if


1) P(b<Z<2.133)=0.1618


1
Expert's answer
2021-07-19T05:45:18-0400
"P(b<Z<2.133)=P(Z<2.133)-P(Z\leq b)"

"\approx0.9835-P(Z\leq b)=0.1618"

"P(Z\leq b)\approx0.8217"

"b\approx0.922"